hello my name is anthony and i am haveing trouble with basic algebra equations and here is what i need help on


When kim works 40 hours in a week,she earns $560. IF she only works 35 hours, how much should she expect to earn?

Found 2 solutions by user_dude2008, Alan3354:
Answer by user_dude2008(1825)

40/560=35/x

40x=35*560

40x=19600

x=19600/40

x=490

Answer: she'll earn $490

Answer by Alan3354(21583)

When kim works 40 hours in a week,she earns $560. IF she only works 35 hours, how much should she expect to earn?
-----------------
$560 for 40 hours = $14 per hour
$14*35 = $490